# TV homogenization inequalities

### v2

@@ 2 Proofs
We will prove a more general, symmetric bound, which immediately implies the one in Theorem 1.1. The constant $2C_{\mathrm{BCV}}$ in the latter’s bound could be sharpened by a factor of $2$ via a direct (asymmetric) proof. L. Mattner [12] has brought to our attention the result of [21], which upper bounds $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ by
@@
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $\frac{2C_{\mathrm{BCV}}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
@@
$\operatorname{TV}(Z,Z+1)=\max_{k}h(k),$ (5)
by a simple telescoping argument. [^3] M. Lutz [12] informs us that this result is contained in [2] with a slightly worse constant. Anti-concentration for Poisson binomials (Baillon–Cominetti–Vaisman). If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [1, Theorem 1] proves the sharp bound
$\max_{k}\mathbb{P}(Z=k)\leq\frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\qquad(v>0),$ (6)
where $\eta_{\mathrm{BCV}}$ is a universal constant ($\eta_{\mathrm{BCV}}\approx 0.4688223555$, [1, Remark 1]). In particular, since $\max_{k}\mathbb{P}(Z=k)\leq 1$ always, we have for all $v\geq 0$
$\max_{k}\mathbb{P}(Z=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{v}}\right\}.$ (7)
Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Edrei-Schoenberg-Whitney theorem [20, Section 4] that the Poisson binomial
270]
has a log-concave, and hence unimodal, law. To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
@@
$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{\operatorname{Var}(T_{i}(t))}}\right\}.$
@@
$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{\bigl(\operatorname{Var}(S(t))-\frac{1}{4}\bigr)_{+}}}\right\}.$
We now upper bound this “singular” expression by a smooth envelope. Set $C_{\mathrm{BCV}}:=\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}$. Then for every $x\geq 0$ we have
$\min\!\left\{1,\ \frac{\eta_{\mathrm{BCV}}}{\sqrt{(x-\frac{1}{4})_{+}}}\right\}\leq\frac{C_{\mathrm{BCV}}}{\sqrt{x+1}}.$ (8)
Indeed, if $x\leq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, the left-hand side equals $1$ and the right-hand side is $C_{\mathrm{BCV}}/\sqrt{x+1}\geq C_{\mathrm{BCV}}/\sqrt{\tfrac{5}{4}+\eta_{\mathrm{BCV}}^{2}}=1$. If $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$, then the left-hand side equals $\eta_{\mathrm{BCV}}/\sqrt{x-\tfrac{1}{4}}$ and squaring shows that (8) is equivalent to $x\geq\tfrac{1}{4}+\eta_{\mathrm{BCV}}^{2}$. Applying (8) with $x=\operatorname{Var}(S(t))$ gives
$\operatorname{TV}\!\bigl(T_{i}(t)+1,T_{i}(t)\bigr)\leq\frac{C_{\mathrm{BCV}}}{\sqrt{\operatorname{Var}(S(t))+1}},$
@@
$|f_{A}^{\prime}(t)|\leq\frac{C_{\mathrm{BCV}}\,\Delta}{\sqrt{\operatorname{Var}(S(t))+1}}.$ (9)
@@
$|f_{A}^{\prime}(t)|\leq\frac{C_{\mathrm{BCV}}\,\Delta}{\sqrt{(1-t)\sigma_{\textup{{\bf q}}}^{2}+t\sigma_{\textup{{\bf p}}}^{2}+1}}.$
@@
$\bigl|f_{A}(1)-f_{A}(0)\bigr|\leq C_{\mathrm{BCV}}\,\Delta\int_{0}^{1}\frac{\mathrm{d}t}{\sqrt{(1-t)(\sigma_{\textup{{\bf q}}}^{2}+1)+t(\sigma_{\textup{{\bf p}}}^{2}+1)}}.$
@@
$\bigl|\mathbb{P}(S_{\textup{{\bf p}}}\in A)-\mathbb{P}(S_{\textup{{\bf q}}}\in A)\bigr|\leq\frac{2C_{\mathrm{BCV}}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
Taking $\sup_{A}$ proves (2.1). ∎
2.2 Proof of Theorem 1.2
We make use of a concentration-variance inequality [13, Disp.(18)], which immediately implies the following “pigeonhole” lemma: [^4] We thank M. Lutz [12] for bringing this to our attention and improving our original lower bound by a factor of $4/3$.
@@
Proof (sketch).
Normalize $g$ to be a probability measure $P$ via $P(\left\{k\right\})=g(k)/G$ and let $X$ be distributed according to $P$. Using the variational characterization of variance, $Var(X)\leq J/G$.
By Mattner–Schulz [13, Disp. (18)], for any law with variance $\sigma^{2}$ and any $h>0$,
@@
Apply this with $h=1$ and $\sigma^{2}=\operatorname{Var}(X)$. Since $P$ is supported on $\mathbb{Z}$, any interval of length $1$ contains at most one atom, and hence
@@
$\frac{1}{G}\max_{k}g(k)\;\geq\;\frac{1}{\sqrt{1+12\operatorname{Var}(X)}}.$
Using $\operatorname{Var}(X)\leq J/G$,
$\max_{k}g(k)\;\geq\;\frac{G}{\sqrt{1+12(J/G)}}=\frac{G^{3/2}}{\sqrt{G+12J}}.$
Finally, since $\sqrt{G+12J}\leq\sqrt{G}+4\sqrt{J}$, we conclude
$\max_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{\,4\sqrt{J}+\sqrt{G}\,}$,
which is the claimed bound. ∎
@@
$g(k):=\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k),\qquad k\in\mathbb{Z}.$ (10)
@@
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;\sup_{k\in\mathbb{Z}}g(k).$ (11)
@@
$J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$ (12)
@@
$\operatorname{Var}_{t}(T_{i}(t))\leq\operatorname{Var}_{t}(S(t))\leq\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}.$ (13)
@@
$\operatorname{TV}(P\times R,\;Q\times S)\leq\operatorname{TV}(P,Q)+\operatorname{TV}(R,S),$ (14)
@@
Acknowledgements
@@
Appendix A Homogenization is a Markov Kernel in the symmetric case
We show that for symmetric $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ — i.e., those satisfying $\textup{{\bf p}}=\textup{{\bf 1}}-\textup{{\bf q}}$, homogenization can be realized by a Markov kernel. We first consider the case $n=2$:
Lemma A.1 (averaging centered Bernoullis).
Let $\alpha,\beta\in(0,1)$ and put $\gamma=(\alpha+\beta)/2$. Then there exists a randomized function $F_{\alpha,\beta}:\left\{0,1\right\}^{2}\to\left\{0,1\right\}^{2}$ with the following property:
$X\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(X)\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),$
$Y\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\alpha}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\beta}{2}\right)$ $\implies$ $F_{\alpha,\beta}(Y)\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right).$
Proof.
As in the proof of [9, Lemma 2.3], it suffices to write down the stochastic vectors representing $X,Y$ and $\hat{X}\sim\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}+\frac{\gamma}{2}\right),\hat{Y}\sim\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)\otimes\operatorname{Ber}\left(\frac{1}{2}-\frac{\gamma}{2}\right)$, the stochastic matrix $M$ representing $F_{\alpha,\beta}$, and to verify that $\hat{X}=XM$ and $\hat{Y}=YM$. For readability, we mildly abuse notation and blur the distinction between the random variable $Z\sim\operatorname{Ber}(p)\otimes\operatorname{Ber}(q)$ and its stochastic vector:
$Z=\left(\mathbb{P}(Z=00),\mathbb{P}(Z=01),\mathbb{P}(Z=10),\mathbb{P}(Z=11)\right)$ $=$ $\left((1-p)(1-q),(1-p)q,p(1-q),pq\right).$
With this convention, the four stochastic vectors are:
$X$ $=$ $\left(\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right)\right),$
$Y$ $=$ $\left(\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}+\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}+\frac{\beta}{2}\right),\left(\frac{1}{2}-\frac{\alpha}{2}\right)\left(\frac{1}{2}-\frac{\beta}{2}\right)\right),$
$\hat{X}$ $=$ $\left(\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right)\right),$
$\hat{Y}$ $=$ $\left(\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}+\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}+\frac{\gamma}{2}\right),\left(\frac{1}{2}-\frac{\gamma}{2}\right)\left(\frac{1}{2}-\frac{\gamma}{2}\right)\right).$
With $\eta:=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}$, we define the $4\times 4$ row-stochastic matrix $M$:
$M=\left(\begin{array}[]{cccc}1&0&0&0\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ \eta&\frac{1}{2}-\eta&\frac{1}{2}-\eta&\eta\\ 0&0&0&1\end{array}\right).$ (19)
Verifying the identities $\hat{X}=XM$ and $\hat{Y}=YM$ is tedious but quite straightforward. Obviously, the rows of $M$ sum to $1$. It only remains to show that the entries of $M$ are nonnegative, or equivalently, that $\eta=\frac{(\alpha-\beta)^{2}}{8(1-\alpha\beta)}\leq 1/2$. The latter is equivalent to ${\alpha^{2}-2\alpha\beta+\beta^{2}}\leq{4(1-\alpha\beta)}$. Expanding and collecting like terms yields the equivalent claim $4-(\alpha+\beta)^{2}\geq 0$, which obviously holds for $\alpha,\beta\in(0,1)$. ∎
For general $n$, the pairwise averaging kernel $M=M(\alpha_{i},\beta_{i})$ defined in (19) can be applied repeatedly to the pairs $(i,j)$. Although the kernel itself changes with each iteration, a standard compactness argument shows that for any symmetric pair $\textup{{\bf p}},\textup{{\bf q}}\in[0,1]^{n}$ there exists a Markov kernel $M_{\textup{{\bf p}},\textup{{\bf q}}}$ simultaneously achieving $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf p}})\mapsto\operatorname{Ber}(\bar{p})^{\otimes n}$ and $M_{\textup{{\bf p}},\textup{{\bf q}}}:\operatorname{Ber}(\textup{{\bf q}})\mapsto\operatorname{Ber}(\bar{q})^{\otimes n}$.

### v3

@@ 2. Proofs
We will prove a more general, symmetric bound, which immediately implies the one in Theorem 1.1. The constant $2C_{0}$ in the latter’s bound could be sharpened by a factor of $2$ via a direct (asymmetric) proof.
@@
$\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})$ $\leq$ $\frac{2C_{0}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
@@
(5) $\operatorname{TV}(Z,Z+1)=\max_{k}h(k),$
by a simple telescoping argument.
Anti-concentration for Poisson binomials (Barbour-Jensen; Baillon–Cominetti–Vaisman). If $Z=\sum_{j=1}^{m}\operatorname{Ber}(r_{j})$ has variance $v=\operatorname{Var}(Z)$, then [2, Theorem 1] proves the sharp bound
(6) $\max_{k}\mathbb{P}(Z=k)\leq\frac{\eta_{0}}{\sqrt{v}}\qquad(v>0),$
where $\eta_{0}$ is a universal constant ($\eta_{0}\approx 0.4688223555$, [2, Remark 1]). [^3] L. Mattner [14] informs us that this result is contained in [3] with a slightly worse constant.
In particular, since $\max_{k}\mathbb{P}(Z=k)\leq 1$ always, we have for all $v\geq 0$
(7) $\max_{k}\mathbb{P}(Z=k)\leq\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{v}}\right\}.$
Unimodality of Poisson binomial pmfs. It is a classic consequence of the Aissen-Edrei-Schoenberg-Whitney theorem [22, Section 4] that the Poisson binomial has a log-concave, and hence unimodal, law. To combine these three ingredients, we begin with the interpolation. For $t\in[0,1]$ define
@@
$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{\operatorname{Var}(T_{i}(t))}}\right\}.$
@@
$\max_{k}\mathbb{P}(T_{i}(t)=k)\leq\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{\bigl(\operatorname{Var}(S(t))-\frac{1}{4}\bigr)_{+}}}\right\}.$
We now upper bound this “singular” expression by a smooth envelope. Set $C_{0}:=\sqrt{\tfrac{5}{4}+\eta_{0}^{2}}$. Then for every $x\geq 0$ we have
(8) $\min\!\left\{1,\ \frac{\eta_{0}}{\sqrt{(x-\frac{1}{4})_{+}}}\right\}\leq\frac{C_{0}}{\sqrt{x+1}}.$
Indeed, if $x\leq\tfrac{1}{4}+\eta_{0}^{2}$, the left-hand side equals $1$ and the right-hand side is $C_{0}/\sqrt{x+1}\geq C_{0}/\sqrt{\tfrac{5}{4}+\eta_{0}^{2}}=1$. If $x\geq\tfrac{1}{4}+\eta_{0}^{2}$, then the left-hand side equals $\eta_{0}/\sqrt{x-\tfrac{1}{4}}$ and squaring shows that (8) is equivalent to $x\geq\tfrac{1}{4}+\eta_{0}^{2}$. Applying (8) with $x=\operatorname{Var}(S(t))$ gives
$\operatorname{TV}\!\bigl(T_{i}(t)+1,T_{i}(t)\bigr)\leq\frac{C_{0}}{\sqrt{\operatorname{Var}(S(t))+1}},$
@@
(9) $|f_{A}^{\prime}(t)|\leq\frac{C_{0}\,\Delta}{\sqrt{\operatorname{Var}(S(t))+1}}.$
@@
$|f_{A}^{\prime}(t)|\leq\frac{C_{0}\,\Delta}{\sqrt{(1-t)\sigma_{\textup{{\bf q}}}^{2}+t\sigma_{\textup{{\bf p}}}^{2}+1}}.$
@@
$\bigl|f_{A}(1)-f_{A}(0)\bigr|\leq C_{0}\,\Delta\int_{0}^{1}\frac{\mathrm{d}t}{\sqrt{(1-t)(\sigma_{\textup{{\bf q}}}^{2}+1)+t(\sigma_{\textup{{\bf p}}}^{2}+1)}}.$
@@
$\bigl|\mathbb{P}(S_{\textup{{\bf p}}}\in A)-\mathbb{P}(S_{\textup{{\bf q}}}\in A)\bigr|\leq\frac{2C_{0}\,\Delta}{\sqrt{\sigma_{\textup{{\bf p}}}^{2}+1}+\sqrt{\sigma_{\textup{{\bf q}}}^{2}+1}}.$
Taking $\sup_{A}$ proves the claim. ∎
2.2. Proof of Theorem 1.2
We make use of a concentration-variance inequality stated in [15, Disp. (18)] and proved therein, which immediately implies the following second-moment extraction lemma. We thank L. Mattner [14] for bringing this to our attention and improving our original lower bound by a factor of $4/3$. We also thank him for the attributional note, “This is due to Paul Lévy in a sharper version, and was in essentially the present version apparently independently discovered by [5, Corollary 2.2]”.
@@
Proof.
Normalize $g$ to be a probability measure $P$ via $P(\left\{k\right\})=g(k)/G$ and let $X$ be distributed according to $P$; then $Var(X)\leq J/G$. By Lévy’s extraction principle, [15, Disp. (18)], for any law with variance $\sigma^{2}$ and any $h>0$,
@@
Apply this with $h=1$ and $\sigma^{2}=\operatorname{Var}(X)$. Since $P$ is supported on $\mathbb{Z}$,
@@
$\max_{k}g(k)$ $\geq$ $\frac{G}{\sqrt{1+12\operatorname{Var}(X)}}\geq\frac{G}{\sqrt{1+12(J/G)}}=\frac{G^{3/2}}{\sqrt{G+12J}}.$
Finally, since $\sqrt{G+12J}\leq\sqrt{G}+4\sqrt{J}$, we conclude $\max_{k\in\mathbb{Z}}g(k)\;\geq\;\frac{G^{3/2}}{\,4\sqrt{J}+\sqrt{G}\,}$, as claimed. ∎
@@
(10) $g(k):=\mathbb{P}(X\geq k)-\mathbb{P}(Y\geq k),\qquad k\in\mathbb{Z}.$
@@
(11) $\operatorname{TV}(S_{\textup{{\bf p}}},S_{\textup{{\bf q}}})\;\geq\;\sup_{k\in\mathbb{Z}}g(k).$
@@
(12) $J=\int_{0}^{1}\sum_{i=1}^{n}\Delta_{i}\,\mathbb{E}_{t}\bigl[(T_{i}(t)+1-\mu_{\textup{{\bf p}}})^{2}\bigr]\,\mathrm{d}t.$
@@
(13) $\operatorname{Var}_{t}(T_{i}(t))\leq\operatorname{Var}_{t}(S(t))\leq\sigma_{\textup{{\bf p}}}^{2}+1+\Delta^{2}.$
@@
(14) $\operatorname{TV}(P\times R,\;Q\times S)\leq\operatorname{TV}(P,Q)+\operatorname{TV}(R,S),$
@@
Appendix A Majorization kernels for complementary Bernoulli products
Fix $n\geq 2$ and write $\Omega:=\{0,1\}^{n}$. For a parameter vector $\alpha=(\alpha_{1},\dots,\alpha_{n})\in[0,1]^{n}$ and $s\in\{+,-\}$ define the product measures
$\mu_{\alpha}^{\,s}:=\bigotimes_{i=1}^{n}\operatorname{Ber}\!\left(\frac{1}{2}+s\,\frac{\alpha_{i}}{2}\right).$
Thus $\mu_{\alpha}^{\,+}$ and $\mu_{\alpha}^{\,-}$ are complementary in the sense that $\mu_{\alpha}^{\,-}$ is obtained from $\mu_{\alpha}^{\,+}$ by flipping each bit.
We show that, in this complementary setting, not only homogenization but any majorization of the parameter vector can be realized by a Markov kernel. We write $\beta\prec\alpha$ to denote that $\beta$ is majorized by $\alpha$, i.e., letting $\alpha^{\downarrow}$ and $\beta^{\downarrow}$ be the vectors obtained by sorting the coordinates of $\alpha$ and $\beta$ in nonincreasing order, we have
$\sum_{i=1}^{k}\beta_{i}^{\downarrow}\ \leq\ \sum_{i=1}^{k}\alpha_{i}^{\downarrow}\quad\text{for all }k=1,\dots,n-1,\qquad\text{and}\qquad\sum_{i=1}^{n}\beta_{i}^{\downarrow}\ =\ \sum_{i=1}^{n}\alpha_{i}^{\downarrow}.$
(see [1, Ch. 1]). We thank L. Mattner [14] for this suggestion; our original result had proven Lemma A.1 only for $\lambda=\frac{1}{2}$ and Theorem A.2 only for homogenization via a compactness argument.
A two–coordinate kernel for a general TT–transform
The basic step is a Markov kernel acting on two coordinates which implements the standard TT–transform (a convex combination of the identity and a transposition).
Lemma A.1 (Two–coordinate TT–transform kernel).
Fix $a,b\in[0,1]$ and $\lambda\in[0,1]$, and define
$a^{\prime}:=\lambda a+(1-\lambda)b,\qquad b^{\prime}:=\lambda b+(1-\lambda)a.$
Let
$\eta:=\begin{cases}\displaystyle\frac{\lambda(1-\lambda)(a-b)^{2}}{2(1-ab)},&ab<1,\\ [1ex]0,&ab=1,\end{cases}$
and consider the row–stochastic $4\times 4$ matrix (indexed by $00,01,10,11$)
(15) $M_{\lambda}(a,b):=\left(\begin{array}[]{cccc}1&0&0&0\\ \eta&\lambda-\eta&(1-\lambda)-\eta&\eta\\ \eta&(1-\lambda)-\eta&\lambda-\eta&\eta\\ 0&0&0&1\end{array}\right).$
Then all entries of $M_{\lambda}(a,b)$ are nonnegative, and for both $s\in\{+,-\}$,
$\Bigl(\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{a}{2}\right)\otimes\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{b}{2}\right)\Bigr)\,M_{\lambda}(a,b)=\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{a^{\prime}}{2}\right)\otimes\operatorname{Ber}\!\left(\tfrac{1}{2}+s\,\tfrac{b^{\prime}}{2}\right).$
Proof.
Row–stochasticity is immediate from (15). For nonnegativity, note that for $a,b\in[0,1]$,
$2(1-ab)-(a-b)^{2}=2-a^{2}-b^{2}\geq 0,$
hence $(a-b)^{2}\leq 2(1-ab)$ whenever $ab<1$. Therefore $\eta\leq\lambda(1-\lambda)\leq\min\{\lambda,1-\lambda\}$, which implies $\lambda-\eta\geq 0$ and $(1-\lambda)-\eta\geq 0$.
To verify the mapping property, write (for $s\in\{+,-\}$) $p_{s}:=\tfrac{1}{2}+s\tfrac{a}{2}$, $q_{s}:=\tfrac{1}{2}+s\tfrac{b}{2}$ and $p_{s}^{\prime}:=\tfrac{1}{2}+s\tfrac{a^{\prime}}{2}$, $q_{s}^{\prime}:=\tfrac{1}{2}+s\tfrac{b^{\prime}}{2}$. Let $x_{s}$ be the stochastic row vector of the input law $\operatorname{Ber}(p_{s})\otimes\operatorname{Ber}(q_{s})$ ordered as $(00,01,10,11)$, and similarly let $y_{s}$ be that of $\operatorname{Ber}(p_{s}^{\prime})\otimes\operatorname{Ber}(q_{s}^{\prime})$. A direct (routine) calculation using (15) shows that $x_{s}\,M_{\lambda}(a,b)=y_{s}$ for $s=+$ and for $s=$-. ∎
Finite decomposition via majorization
Lemma A.1 supplies the probabilistic implementation of a single TT–transform. The following theorem is the desired global statement.
Theorem A.2 (Majorization kernel for complementary products).
Fix $n\geq 2$ and $\alpha,\beta\in[0,1]^{n}$ with $\beta\prec\alpha$. Then there exists a Markov kernel $K:\Omega\to\mathcal{P}(\Omega)$ such that simultaneously, for both $s\in\{+,-\}$,
$\mu_{\alpha}^{\,s}K=\mu_{\beta}^{\,s}.$
Proof.
By the Muirhead–Hardy–Littlewood–Pólya lemma (see [1, Lemma B.1]), there exist indices $(i_{t},j_{t})$ and parameters $\lambda_{t}\in[0,1]$, $t=1,\dots,m$, such that if we define $\alpha^{(0)}:=\alpha$ and for each $t\geq 1$ update only coordinates $(i_{t},j_{t})$ via
$\bigl(\alpha^{(t)}_{i_{t}},\alpha^{(t)}_{j_{t}}\bigr)=\bigl(\lambda_{t}\alpha^{(t-1)}_{i_{t}}+(1-\lambda_{t})\alpha^{(t-1)}_{j_{t}},\ \lambda_{t}\alpha^{(t-1)}_{j_{t}}+(1-\lambda_{t})\alpha^{(t-1)}_{i_{t}}\bigr)$,
with all other coordinates unchanged, then $\alpha^{(m)}=\beta$.
For each step $t$, let $K_{t}$ be the Markov kernel on $\Omega$ that acts only on coordinates $(i_{t},j_{t})$ as follows: given $x\in\Omega$, sample
$(x^{\prime}_{i_{t}},x^{\prime}_{j_{t}})\sim M_{\lambda_{t}}\!\left(\alpha^{(t-1)}_{i_{t}},\alpha^{(t-1)}_{j_{t}}\right)(x_{i_{t}},x_{j_{t}})$
(using Lemma A.1), and set $x^{\prime}_{k}=x_{k}$ for $k\neq i_{t},j_{t}$. By independence of coordinates and Lemma A.1, for both $s\in\{+,-\}$ we have
$\mu_{\alpha^{(t-1)}}^{\,s}K_{t}=\mu_{\alpha^{(t)}}^{\,s}.$
Finally, define the composed kernel $K:=K_{1}K_{2}\cdots K_{m}$. Iterating the last display yields $\mu_{\alpha}^{\,s}K=\mu_{\beta}^{\,s}$ for both $s\in\{+,-\}$. ∎
Corollary A.3 (Homogenization as a special case).
Let $\alpha\in[0,1]^{n}$ and put $\bar{\alpha}:=\frac{1}{n}\sum_{i=1}^{n}\alpha_{i}$. Then there exists a Markov kernel $K$ on $\Omega$ such that for both $s\in\{+,-\}$,
$\mu_{\alpha}^{\,s}K=\mu_{\bar{\alpha}\mathbf{1}}^{\,s}.$
Proof.
It is immediate that $\bar{\alpha}\mathbf{1}\prec\alpha$. Apply Theorem A.2 with $\beta=\bar{\alpha}\mathbf{1}$. ∎
Acknowledgments
@@
This research was supported in part by the Israel Science Foundation ISF grant 581/25 and the Binational Science Foundation BSF grant 2024243.